# Rajamanickam

### v2

@@ 5 Oseen–Frank limit, $\varepsilon\to\infty$
$\mathbf{Q}_{b}=(\mathbf{n}_{b}\otimes\mathbf{n}_{b}-\tfrac{1}{2}\mathbf{I}),\qquad\mathbf{n}_{b}\begin{bmatrix}\cos\varphi_{b}\\ \sin\varphi_{b}\end{bmatrix}$, (30)
@@ 5.1 Outer solution
$\mathbf{Q}=\mathbf{Q}_{0}+\varepsilon^{-1}\varphi_{1}\mathbf{T}_{0}+\varepsilon^{-2}[\varphi_{2}\mathbf{T}_{0}-2(|\nabla\varphi_{0}|^{2}+\varphi_{1}^{2})\mathbf{Q}_{0}]+\cdots$ (44)
@@
$\mathbf{Q}_{0}=\mathbf{n}_{0}\otimes\mathbf{n}_{0}-\tfrac{1}{2}\mathbf{I}_{2},\qquad\mathbf{T}_{0}=\mathbf{n}_{0}\otimes\mathbf{t}_{0}+\mathbf{t}_{0}\otimes\mathbf{n}_{0}$ (45)
and $\mathbf{Q}_{0}:\mathbf{T}_{0}=0$. In the limit $\gamma\to\infty$ (i.e., Dirichlet boundary conditions), $\varphi_{1}=0$. For this case, $O(\varepsilon^{-2})$ term proportional to $\mathbf{Q}_{0}$ was obtained earlier in [di2020landau, Section 4] as an $O(\varepsilon^{-2})$ correction affecting only the scalar order parameter. Our expansion provides the full $O(\varepsilon^{-2})$ correction in $\mathbf{Q}$, including the component $\varphi_{2}\mathbf{T}_{0}$ orthogonal to $\mathbf{Q}_{0}$, which corresponds to a second-order correction in the director field not previously computed.
@@ Acknowledgment
Let $(\bar{q}_{1},\bar{q}_{2})$ be a critical point of the free energy functional (10). To examine its stability, we study the associated gradient–flow dynamics
$\frac{\partial q_{1}}{\partial t}=-\frac{\delta F}{\delta q_{1}},\qquad\frac{\partial q_{2}}{\partial t}=-\frac{\delta F}{\delta q_{2}}.$ (76)
@@
where $t$ is a fictitious time that drives the system towards lower energy states. By superposing small perturbations to $(\bar{q}_{1},\bar{q}_{2})$ of the form,
$q_{1}=\bar{q}_{1}(x,y)+\hat{q}_{1}(x,y)e^{-\sigma t},\qquad q_{2}=\bar{q}_{2}(x,y)+\hat{q}_{2}(x,y)e^{-\sigma t}$ (77)
@@
with $|\hat{q}_{1}|\ll|\bar{q}_{1}|$ and $|\hat{q}_{2}|\ll|\bar{q}_{2}|$ and linearising the gradient-flow equations, we obtain the eigenvalue problem [rajamanickam2026nematic]
$\mathbf{H}_{(\bar{q}_{1},\bar{q}_{2})}\begin{bmatrix}\hat{q}_{1}\\ \hat{q}_{2}\end{bmatrix}=\sigma\begin{bmatrix}\hat{q}_{1}\\ \hat{q}_{2}\end{bmatrix}$ (78)
where $\mathbf{H}_{(\bar{q}_{1},\bar{q}_{2})}$ is the Hessian evaluated at the critical point and is given by
@@
$\mathbf{H}_{(\bar{q}_{1},\bar{q}_{2})}=\left.\frac{\delta^{2}F}{\delta q_{1}\delta q_{2}}\right|_{(\bar{q}_{1},\bar{q}_{2})}=2\begin{bmatrix}-\nabla^{2}&0\\ 0&-\nabla^{2}\end{bmatrix}+2\varepsilon^{2}\begin{bmatrix}12\bar{q}_{1}^{2}+4\bar{q}_{2}^{2}-1&8\bar{q}_{1}\bar{q}_{2}\\ 8\bar{q}_{1}\bar{q}_{2}&4\bar{q}_{1}^{2}+12\bar{q}_{2}^{2}-1\end{bmatrix}$. (79)
The eigenvalue problem must be solved using the homogeneous boundary condition
@@
$\frac{\partial\hat{q}_{1}}{\partial\nu}=-\varepsilon\gamma\hat{q}_{1},\quad\frac{\partial\hat{q}_{2}}{\partial\nu}=-\varepsilon\gamma\hat{q}_{2}\quad\text{on}\quad\partial\Omega.$ (80)
@@
Since $\mathbf{H}$ is self-adjoint, all its eigenvalues are real. The spectrum is discrete and may be arranged in non-decreasing order, $\sigma_{1}\leq\sigma_{2}\leq\sigma_{3}\leq\dots$. The critical point $(\bar{q}_{1},\bar{q}_{2})$ is asymptotically stable when $\sigma_{1}>0$, marginally stable when $\sigma_{1}=0$ and unstable when $\sigma_{1}<0$. The number of negative eigenvalues gives the Morse index of the equilibrium and quantifies the number of linearly unstable directions in the energy landscape.

### v3

@@ 4 Small-domain limit, $\varepsilon\to 0$
Previous work by Fang et al. [fang2020surface] considered surface energy effects in the limit $\varepsilon\to 0$, but under the scaling $\gamma\varepsilon\sim 1$, for which a well-ordered structure is obtained at leading order. While it is noted there that $\gamma\varepsilon$ must vanish as $\varepsilon\to 0$, the leading-order solution is taken to be an arbitrary constant. By contrast, in the present analysis, the leading-order state is uniquely fixed by the boundary anchoring through the solvability condition.
@@ 5 Oseen–Frank limit, $\varepsilon\to\infty$
$\mathbf{Q}_{b}=(\mathbf{n}_{b}\otimes\mathbf{n}_{b}-\tfrac{1}{2}\mathbf{I}),\qquad\mathbf{n}_{b}=\begin{bmatrix}\cos\varphi_{b}\\ \sin\varphi_{b}\end{bmatrix}$, (30)
@@ 5.1 Outer solution
$\mathbf{Q}=\mathbf{Q}_{0}+\varepsilon^{-1}\varphi_{1}\mathbf{T}_{0}+\varepsilon^{-2}[\varphi_{2}\mathbf{T}_{0}-(|\nabla\mathbf{Q}_{0}|^{2}+2\varphi_{1}^{2})\mathbf{Q}_{0}]+\cdots$ (44)
@@
$\mathbf{Q}_{0}$ $=\left.\frac{\partial\mathbf{Q}}{\partial s}\right|_{s=1,\varphi=\varphi_{0}}=\mathbf{n}_{0}\otimes\mathbf{n}_{0}-\tfrac{1}{2}\mathbf{I}_{2},$ (45)
$\mathbf{T}_{0}$ $=\left.\frac{\partial\mathbf{Q}}{\partial\varphi}\right|_{s=1,\varphi=\varphi_{0}}=\mathbf{n}_{0}\otimes\mathbf{t}_{0}+\mathbf{t}_{0}\otimes\mathbf{n}_{0}$ (46)
with $\mathbf{Q}_{0}:\mathbf{T}_{0}=0$. In the limit $\gamma\to\infty$ (i.e., Dirichlet boundary conditions), $\varphi_{1}=0$. For this case, $O(\varepsilon^{-2})$ term proportional to $\mathbf{Q}_{0}$ was obtained earlier by Di Fratta et. al. [di2020landau, Section 4] as an $O(\varepsilon^{-2})$ correction affecting only the scalar order parameter. Our expansion provides the full $O(\varepsilon^{-2})$ correction in $\mathbf{Q}$, including the component $\varphi_{2}\mathbf{T}_{0}$ orthogonal to $\mathbf{Q}_{0}$, which corresponds to a second-order correction in the director field not previously computed.
@@ Acknowledgment
Let $\bar{\mathbf{q}}=(\bar{q}_{1},\bar{q}_{2})^{T}$ be a critical point of the free energy functional $F[\mathbf{Q}]$ gievn in (10). To examine its stability, we study the associated gradient–flow dynamics
$\frac{\partial\mathbf{q}}{\partial t}=-\frac{\delta F}{\delta\mathbf{q}},\quad\text{where}\quad\mathbf{q}=\begin{bmatrix}q_{1}\\ q_{2}\end{bmatrix}$ (77)
@@
and $t$ is a fictitious time that drives the system towards lower energy states. By superposing small perturbations to $\bar{\mathbf{q}}$ of the form,
$\mathbf{q}=\bar{\mathbf{q}}(x,y)+\hat{\mathbf{q}}(x,y)e^{-\sigma t},$ (78)
@@
with $|\hat{\mathbf{q}}|\ll|\bar{\mathbf{q}}|$ and linearising the gradient-flow equations, we obtain the eigenvalue problem [rajamanickam2026nematic]
$\mathbf{H}_{\bar{\mathbf{q}}}\mathbf{q}=\sigma\mathbf{q}$ (79)
where $\mathbf{H}_{\bar{\mathbf{q}}}$ is the Hessian evaluated at the critical point and is given by
@@
$\mathbf{H}_{\bar{\mathbf{q}}}=\left.\frac{\delta^{2}F}{\delta q_{1}\delta q_{2}}\right|_{\bar{\mathbf{q}}}=2\begin{bmatrix}-\nabla^{2}&0\\ 0&-\nabla^{2}\end{bmatrix}+2\varepsilon^{2}\begin{bmatrix}12\bar{q}_{1}^{2}+4\bar{q}_{2}^{2}-1&8\bar{q}_{1}\bar{q}_{2}\\ 8\bar{q}_{1}\bar{q}_{2}&4\bar{q}_{1}^{2}+12\bar{q}_{2}^{2}-1\end{bmatrix}$. (80)
The perturbations are required to satisfy the homogeneous boundary condition
@@
$\frac{\partial\hat{\mathbf{q}}}{\partial\nu}=-\varepsilon\gamma\hat{\mathbf{q}}\quad\text{on}\quad\partial\Omega.$ (81)
@@
Since $\mathbf{H}$ is self-adjoint, all its eigenvalues are real. The spectrum is discrete and may be arranged in non-decreasing order, $\sigma_{1}\leq\sigma_{2}\leq\sigma_{3}\leq\dots$. The critical point $\bar{\mathbf{q}}$ is asymptotically stable when $\sigma_{1}>0$, marginally stable when $\sigma_{1}=0$ and unstable when $\sigma_{1}<0$. The number of negative eigenvalues gives the Morse index of the equilibrium and quantifies the number of linearly unstable directions in the energy landscape.
Stability in the Oseen–Frank limit
It is instructive to deduce the formulation of the stability analysis from (79) for the leading-order Oseen–Frank problem described in (35)-(36). For shortness, we shall drop the subscript ”0” for the leading-order solution. First of all, let us substitute $\bar{s}^{2}=4\bar{q}^{2}=1$, so that the eigenvalue equation reduces to
$\nabla^{2}\hat{\mathbf{q}}-8\varepsilon^{2}\bar{\mathbf{q}}\bar{\mathbf{q}}^{T}\hat{\mathbf{q}}=-\frac{\sigma}{2}\hat{\mathbf{q}}$. (82)
The matrix $\bar{\mathbf{q}}\bar{\mathbf{q}}^{T}$ is of rank one, with eigenvalues $(\frac{1}{4},0)$ and corresponding eigenvectors $(\mathbf{e}_{\parallel}=\bar{\mathbf{q}}=(\bar{q}_{1},\bar{q}_{2})^{T},\mathbf{e}_{\perp}=(-\bar{q}_{2},\bar{q}_{1})^{T})$. Hence any perturbation can be decomposed as
$\hat{\mathbf{q}}=\hat{q}_{{}_{\parallel}}\mathbf{e}_{\parallel}+\hat{q}_{\perp}\mathbf{e}_{\perp}.$ (83)
However, in the Oseen–Frank limit, $\varepsilon\to\infty$, the linear perturbation must remain on the manifold $s=1$, so that only the perpendicular component is dynamically relevant. The resulting eigenvalue equation reads
$\nabla^{2}\hat{q}_{\perp}+(\tfrac{1}{2}\sigma-4|\nabla\bar{\varphi}|^{2})\hat{q}_{\perp}=0$ (84)
where the term $4|\nabla\bar{\varphi}|^{2}\hat{q}_{\perp}$ arises from the spatial variation of the basis vector $\mathbf{e}_{\perp}(\mathbf{x})$. Introducing $\hat{q}_{\perp}=2\hat{\varphi}$, with $\hat{\varphi}$ denoting the perturbation of the director field, one obtains
$\nabla^{2}\hat{\varphi}+(\tfrac{1}{2}\sigma-4|\nabla\bar{\varphi}|^{2})\hat{\varphi}=0$ (85)
subject to $\hat{\varphi}=0$ on $\partial\Omega$. This eigenvalue problem is formally analogous to the two-dimensional Schrödinger equation
$\nabla^{2}\psi+\frac{2m}{\hbar^{2}}(E-V)\psi=0$ (86)
under the correspondence
$\hat{\varphi}\longmapsto\psi,\qquad 1\longmapsto\frac{2m}{\hbar^{2}},\qquad\frac{\sigma}{2}\longmapsto E,\qquad 4|\nabla\bar{\varphi}|^{2}\longmapsto V.$ (87)
Since the potential $4|\nabla\bar{\varphi}|^{2}\geq 0$, it follows that the energy $\sigma/2\geq 0$, indicating that the base state is linearly stable. This reflects the strength and limitation of the Oseen–Frank approximation because elastic stability of the director field is captured while instabilities associated with variations in the magnitude of $\mathbf{Q}$ are neglected. In the full Landau–de Gennes theory, variations in $s$ can produce negative eigenvalues corresponding to core instabilities of defects. Within such cores, both $\hat{q}_{\parallel}$ and $\hat{q}_{\perp}$ must be retained, and the above analysis provides only an approximate description.